# Robust $\mathcal{H}_{\infty}$ control and worst-case search in constrained parametric space

### v4

@@ Abstract
Standard $\mathcal{H}_{\infty}$$\mathcal{H}_{2}$ robust control and analysis tools operate on uncertain parameters assumed to vary independently within prescribed bounds. This paper extends their capabilities in the presence of constraints coupling these parameters and restricting the parametric space. Focusing on the worst-case search, we demonstrate – based on the theory of upper-$C^{1}$ functions – the validity of using standard, readily available smooth optimization algorithms to address this nonsmooth constrained optimization problem. In particular, we prove that the sequential quadratic programming algorithm converges to Karush-Kuhn-Tucker points, and that such conditions are satisfied by any subgradient at a local minimum. This worst-case search then enables robust controller synthesis: identified worst-case configurations are iteratively added to an active set on which a non-smooth multi-models optimization of the controller is performed. The methodology is illustrated on a satellite benchmark with flexible appendages, of order 50 with 43 uncertain parameters. From a practical point of view, we combine the local exploitation proposed above with a global exploration using either Monte-Carlo sampling or particle swarm optimization. We show that the proposed constrained optimization effectively complements Monte-Carlo sampling by enabling fast detection of rare worst-case configurations, and that the robust controller optimization converges with less than 10 active configurations.
@@ I Introduction
The standard robust structured $\mathcal{H}_{\infty}$ control problem [10] is formulated as the min-max optimization problem:
@@
where $T_{zw}(\mathrm{s},K,\delta)$ is a transfer function between input $w$ and output $z$, $K$ is the controller to be optimized, $\mathcal{K}$ is the set of controllers of the chosen structure, and $\delta$ is the vector of $k$ uncertain parameters in a parametric space $\mathcal{D}$. A multi-objective formulation can also be stated as follows:
@@
where $||\cdot||$ may refer to the $\mathcal{H}_{2}$ or $\mathcal{H}_{\infty}$ system norm. The channel $w_{1}\rightarrow z_{1}$ corresponds to a ”soft” requirement to be minimized, while the other channels $w_{i}\rightarrow z_{i}$ ($i>1$) are ”hard” requirements, that must be inferior to 1 (once normalized) but not necessarily minimized.
In this paper, we mainly focus on the inner maximization problems, which will then allow us to solve Problems (1) and (2). Namely, search for worst-case performance:
@@
where $A(K,\delta)$ is the matrix A of the state-space representation of the closed-loop system, and $\alpha(\cdot)$ denotes the spectral abscissa:
@@
Classically in the literature, $\mathcal{D}$ is the hypercube $1,1]^{k}$ after normalization of the uncertain parameters, and $T_{zw}(s,K,\delta)$ is modeled under the Linear Fractional Representation (LFR). In this case, MATLAB’s function systune, relying on the non-smooth optimization of [3, 8], is now widely used in academia and industry to solve problems (1) and (2). More specifically, in [3], problems (3) and (4) are solved with local non-smooth optimization. Indeed, the computation of Clarke subdifferentials [22] of the spectral abscissa and of the $\mathcal{H}_{\infty}$ norm [16] enables a bundle method to quickly find worst-case configurations with regard to stability and performance. The control problem (1) is then tackled by combining this worst-case search with the non-smooth controller optimization of [8]. This approach relies on the construction of an inner approximation $\mathcal{D}_{a}\subset\mathcal{D}$ composed of so-called active configurations. At each iteration, the controller is tuned so as to be robust to all current active configurations. Then, a worst-case search is performed; if the worst-case configuration does not satisfy the control requirements, it is added to $\mathcal{D}_{a}$ and the algorithm goes back to controller tuning; otherwise, the algorithm can terminate. It is shown that this approach is both more reliable and more efficient than gridding the parametric space, which quickly becomes intractable when the number of parameters increases. In practice, the optimization generally finishes with a small number of active configurations. As this approach relies on local searches, robustness is generally thoroughly verified a posteriori. $\mu$-analysis [29] was developed for this purpose, generally based on branch-and-bound algorithms [47]. It provides guaranteed deterministic lower and upper bounds of the global optimum; a probabilistic variant also exists [13, 50]. Monte-Carlo sampling can also guarantee a probability of robustness within predefined thresholds of accuracy and confidence [53]. Another common industrial practice is to grid the parametric space with a limited number of samples, typically 100 to 1000 combinations of the parameters bounds. But this approach will miss any worst case occurring between grid points and, unlike Monte-Carlo, does not provide any statistical guarantee.
@@
Our main focus is to seek computationally efficient techniques that can be integrated into robust control design, alternating between multi-models controller synthesis and worst-case search as in [3], to solve the standard problem (1) and its multi-objectives formulation (2) in reasonable time.
The robust control and analysis approaches discussed above do not address the constrained case, at least not in their current implementations. Nevertheless, quantifying robustness can still be approached with Monte-Carlo sampling; samples that do not satisfy the nonlinear constraint are simply discarded. Evaluating the stability or the norm of a linear system at a given parametric configuration is relatively fast, and easily parallelizable; moreover, the number of evaluations required to provide the desired levels of probability and confidence is independent from the number of parameters [53]. However, finding a worst case, e.g. to address problems (3) or (4), may require numerous function evaluations and lead to a significant computation time, especially in critical systems for which even rare worst cases must be avoided. This is particularly an issue when the robustness analysis iterates with controller design to perform robust tuning.
@@
In this paper, we demonstrate new properties of the class of upper-$C^{1}$ functions [51], and in particular the convergence of Sequential Quadratic Programming (SQP) [38], a standard constrained optimization algorithm, when applied to such functions even though they are not smooth. We deduce that SQP can be applied to the nonsmooth problem (3), and to (4) under an additional assumption, to find their Karush-Kuhn-Tucker points. From a practical point of view, we also study the use of Particle Swarm Optimization (PSO) [34], tackling the nonlinear constraint through a penalty approach [27], or Monte-Carlo sampling (if stochastic guarantees are sought) to perform a global exploration of the parameter space, with SQP subsequently ensuring efficient local exploitation. Then, following the approach of [3], this worst-case search method is then combined with multi-models controller optimization [8] to perform robust control. To the best of our knowledge, this is the first robust $\mathcal{H}_{\infty}$ control approach that tackles nonlinear constraints on the uncertain parameters.
@@
We emphasize that the proposed techniques, although studied here in the context of the constrained space (5), may also prove valuable in the unconstrained case. Indeed, $\mu$-analysis remains difficult to apply to industrial systems with many uncertain parameters – as illustrated in the recent references [33, 32, 36, 37] where systems with 16 to 20 parameters required sensitivity-based model reduction to make the problem tractable, at the expense of model fidelity. In particular, [49] shows that such $\mu$-analysis can even lead to an overly optimistic worst-case prediction after model reduction. This occurs because the formal robustness guarantees apply only to the reduced model and do not extend to the original system, thus limiting the practical value of sensitivity-based $\mu$-analysis. By contrast, the proposed approach scales well to plants with many states and parameters (as a gradient-based method), and thus can be applied to the full system. Furthermore, the non-smooth optimization of [3], implemented in systune, is local and sometimes misses worst-case configurations, whereas the proposed approach performs a global exploration of the parametric space, albeit without formal guarantees of global optimality.
We illustrate the proposed approach on the robust control of a spacecraft with flexible appendage. In such systems, uncertainties associated with the mechanical properties of flexible appendages must satisfy a condition of the form (5) or otherwise be excluded from robustness analysis (specifically, the residual mass matrix must remain positive definite). While structured $\mathcal{H}_{\infty}$ control has become industrial standard practice among satellite integrators for missions with demanding control requirements [36, 52, 30, 20], this constraint, and more generally the robustness analysis, is generally handled using Monte-Carlo or gridding approaches, which suffer from the aforementioned drawbacks. The proposed methodology offers a practical and scalable alternative that can significantly reduce engineering effort and computational cost.
@@ II-A Properties of upper-$C^{1}$ functions
[3, Sections III.A., II.B., V.A.] shows that minimization of such upper-$C^{1}$ functions, even when not differentiable everywhere, exhibits similar behavior to smooth functions. In particular, the authors show the convergence of their bundle method, in the search of the worst-case $\mathcal{H}_{\infty}$ norm, even when the cutting plane step is reduced to a standard linesearch (as in smooth optimization). This work follows [26] that treats of the optimization of lower- and upper-$C^{1}$ functions. Similar considerations are also thoroughly discussed by the same authors in [11]. Similarly, in [28], the Frank-Wolfe algorithm, originally designed for smooth functions, is also proven to converge to Clarke-stationnary points when applied to upper-$C^{1}$ functions. In the same spirit, we now show that the upper-$C^{1}$ functions, although not differentiable everywhere, maintain some favorable properties for their minimization. Then, we will demonstrate the convergence of one particular smooth optimization algorithm when applied to upper-$C^{1}$ functions.
@@
The directional derivative at $x$ in direction $d$ of $f$ defined as
@@
always exists for an upper-$C^{1}$ function [46].
@@
Let a locally Lipschitz function $f$ be upper-$C^{1}$ at $x$, and $g_{x}\in\partial f(x)$ any subgradient of $f$ at $x$. We consider Problem (6).
@@
b) Optimality. If $x$ (locally) minimizes $f(x)$, then $g_{x}||=0$ for any $g_{x}\in\partial f(x)$.
@@
a) [28, Proposition 2 and equation (6)] shows that the directional derivative of an upper-$C^{1}$ function $f$ is:
$f^{\prime}(x,d)=\min_{g\in\partial f(x)}\langle g,d\rangle=\min_{u\in I(x)}\langle\nabla_{x}F(x,u),d\rangle$ (7)
@@
Let a locally Lipschitz function $f$ be upper-$C^{1}$ at $x$, and $g_{x}\in\partial f(x)$ any subgradient of $f$ at $x$. Let the $c_{i}$ be continuously differentiable. We consider Problem (8).
a) Descent direction. Let $(p,u)$ be a KKT pair of $Q(x,H,g_{x})$. If $p||\neq 0$ and $u||_{\infty}\leq r$, then $p$ is a descent direction for $\theta_{r}(x)$, i.e. $\theta^{\prime}_{r}(x,p)<0$.
b) Optimality. If $x$ (locally) minimizes $f(x)$ subject to $n$ constraints $c_{i}(x)\leq 0$ (Problem (8)), then any $g_{x}\in\partial f(x)$ verifies the following KKT conditions: there exists $(\mu_{i})_{1\leq i\leq n}$ such that:
@@
a) Cf. Appendix. The proof is a mere adaptation of [31, Theorem 3.1] to the upper-$C^{1}$ case.
@@
Proposition 1.b) is similar to [28, Proposition 4], where it is shown that ”directional-stationarity”, i.e. the assumption that $f^{\prime}(x,d)\geq 0$ for any $d\in\mathbb{R}^{k}$, implies that all subgradients are 0. Note that directional stationarity does not imply that $x$ minimizes $f(x)$; for this, more assumptions would be required, e.g. convexity of the $F(\cdot,s)$ [28, Theorem 10]. Proposition 2.a) is a generalization of [31, Theorem 3.1], originally for differentiable functions, to upper-$C^{1}$ functions. To the best of our knowledge, Proposition 2.b) is novel, and can also be seen as a specific case of the usual KKT conditions for upper-$C^{1}$ functions. It allows optimality to be checked with any subgradient, in contrast to the general nonsmooth case, where only the existence of one subgradient satisfying KKT conditions is required.
@@ II-B Convergence of sequential quadratic programming (SQP) for upper-$C^{1}$ functions
We now introduce the Sequential Quadratic Programming (SQP) algorithm to solve Problem (8). It is described in Algorithm 1, as in [31] except that we replaced the gradient $\nabla f(x)$ by any subgradient $g_{x}\in\partial f(x)$ in the quadratic problem $Q(x,H,g_{x})$ (equation (10)). See also [44] for a discussion when the penalty parameter $r$ is variable. We do not detail how the quadratic problem in Step 1 or the step size $\lambda_{k}$ in Step 2 may be solved, or how $H_{k}$ is updated in Step 3, as it does not affect the convergence proof (as long as the conditions described in Algorithm 1 are satisfied)
@@
Step 1: Pick a $g_{x_{k}}\in\partial f(x_{k})$. Find a KKT point $p_{k}$ of the quadratic problem $Q(x_{k},H_{k},g_{x_{k}})$ defined in equation (10).
@@
where $\theta_{r}$ is defined in equation (11), and ($\epsilon_{k}$) is a sequence of nonnegative numbers satisfying:
@@
Proposition 3 extends the convergence result of SQP established in [31, Theorem 3.2] for differentiable objective functions to the case of upper-$C^{1}$ objective functions. Besides the upper-$C^{1}$ condition of $f$, the same assumptions as in [31, Theorem 3.2] are considered.
@@
(ii) For each $k$, there exists a KKT point of $Q(x_{k},H_{k},g_{x_{k}})$ (equation (10)) with a Lagrange multiplier vector $u_{k}$ such that $u_{k}||_{\infty}\leq r$ (vector $\infty$-norm).
@@
Cf. Appendix. The proof follows that of [31, Theorem 3.2] and adapts it to the case where f is not differentiable but only upper-$C^{1}$. In particular, the upper semi-continuity of the Clarke subdifferential allows to maintain the argument linked to the continuity of the gradient in the original proof. The existence of an accumulation point of the solutions $p_{k}$ of the quadratic problems $Q(x_{k},H_{k},g_{x_{k}})$ also follows from the submonotonicity of the subdifferential of lower-$C^{1}$ functions [51] (cf. Lemma 1 in Appendix). ∎
@@
The purpose of our arguments for using smooth optimization is not to claim superiority over specialized nonsmooth optimization methods (for example [26, 25] for lower- or upper-$C^{1}$ functions specifically). Instead, we emphasize that readily available algorithms (e.g., in MATLAB or Python toolboxes) can be applied successfully and produce satisfactory results.

### v5

@@ Abstract
Standard $\mathcal{H}_{\infty}$$\mathcal{H}_{2}$ robust control and analysis tools operate on uncertain parameters assumed to vary independently within prescribed bounds. This paper extends their capabilities in the presence of constraints coupling these parameters and restricting the parametric space. Focusing on the worst-case search, we demonstrate – based on the theory of upper-$C^{1}$ functions – the validity of standard, readily available smooth optimization to address this nonsmooth constrained optimization problem. Specifically, we prove that the sequential quadratic programming algorithm converges to Karush-Kuhn-Tucker points, and that such conditions are satisfied by any subgradient at a local minimum. This worst-case search then enables robust controller synthesis: identified worst-case configurations are iteratively added to an active set on which a non-smooth multi-models optimization of the controller is performed. From a practical point of view, we combine the local exploitation proposed above with a global exploration using Monte-Carlo sampling. The proposed approach is illustrated through the robust control of a mechanical system, of order 50 with 43 uncertain parameters. We show that the constrained optimization effectively complements Monte-Carlo sampling by enabling fast detection of rare worst-case configurations, and that the robust controller optimization converges with less than 10 active configurations.
@@ I Introduction
The $\mathcal{H}_{\infty}$ norm is a widely used metric to quantify the performance of control systems. For a linear system, it is equivalently defined as the $\mathcal{L}_{2}$-induced norm, hence representing the maximum energy amplification from input to output, or as the maximum peak in the frequency domain:
@@
where $T_{zw}(\mathrm{s})$ is a transfer function between input $w$ and output $z$. Another system norm is the $\mathcal{H}_{2}$ norm, representing the energy of the impulse response, or the output’s standard deviation in response to a white noise of unit power spectral density. It is thus natural to seek to optimize these norms in control problems. Moreover, practical problems include parametric uncertainties that may severely degrade the performance and the stability if not addressed during controller optimization. Consequently, the standard robust structured $\mathcal{H}_{\infty}$ or $\mathcal{H}_{2}$ control problem is formulated as the min-max optimization problem:
@@
where $||\cdot||$ may refer to the $\mathcal{H}_{2}$ or $\mathcal{H}_{\infty}$ system norm, and the transfer function $T_{zw}(\mathrm{s},K,\delta)$ is now written as a function of $K$ the controller to be optimized, $\mathcal{K}$ is the set of controllers of the chosen structure, and $\delta$ is the vector of $k$ uncertain parameters in a parametric space $\mathcal{D}$. In presence of multiple control requirements, the following formulation may be used:
[…]
Classically in the literature, $\mathcal{D}$ is the hypercube $1,1]^{k}$ after normalization of the uncertain parameters. In this context, an overview of this control problem and of the modern resolution method is presented in [10]. The goal of this paper is to address the presence of nonlinear constraints restricting this hypercube.
We mainly focus on the inner maximization problems – it will then allow us to solve Problems (1) and (2) by leveraging state-of-art nonsmooth optimization of the controller [8]. Namely, search for worst-case performance:
@@
where $A(K,\delta)$ is the matrix A of the state-space representation of the closed loop, and $\alpha(\cdot)$ denotes the spectral abscissa:
@@
When the parametric space is the whole hypercube $1,1]^{k}$, Matlab’s function systune, relying on the non-smooth optimization of [8, 3], is now widely used in academia and industry to solve problems (1) and (2). More specifically, in [3], problems (3) and (4) are solved with local non-smooth optimization. Indeed, the computation of Clarke subdifferentials [21] of the spectral abscissa and of the $\mathcal{H}_{\infty}$ norm [16] enables a bundle method to quickly find worst-case configurations with regard to stability and performance. The control problem (1) is then tackled by combining this worst-case search with the non-smooth controller optimization of [8]. This approach relies on the construction of an inner approximation $\mathcal{D}_{a}\subset\mathcal{D}$ composed of so-called active configurations. At each iteration, the controller is tuned so as to be robust to all current active configurations. Then, a worst-case search is performed; if the worst-case configuration does not satisfy the control requirements, it is added to $\mathcal{D}_{a}$ and the algorithm goes back to controller tuning; otherwise, the algorithm can terminate. It is shown that this approach is both more reliable and more efficient than gridding the parametric space, which quickly becomes intractable when the number of parameters increases. In practice, the optimization generally finishes with a small number of active configurations. As this approach relies on local searches, robustness is generally thoroughly verified a posteriori. $\mu$-analysis [27] was developed for this purpose, generally based on branch-and-bound algorithms [36]. It provides guaranteed deterministic lower and upper bounds of the global optimum; a probabilistic variant also exists [13, 38]. Monte-Carlo sampling can also guarantee a probability of robustness within predefined thresholds of accuracy and confidence [40]. Another common industrial practice is to grid the parametric space with a limited number of samples. But this approach has exponential complexity with respect to the number of parameters, will miss any worst case occurring between grid points and does not provide any statistical guarantee.
@@
The robust control and analysis approaches discussed above do not address the constrained case, at least not in their current implementations. Our focus is twofold. Firstly, we seek computationally efficient techniques that can be integrated into robust control design, alternating between multi-models controller synthesis and worst-case search as in [3], to solve the standard problem (1) and its multi-objectives formulation (2) in reasonable time. Secondly, we aim to assess their reliability in consistently detecting global worst-case configurations as a validation tool.
@@
Specifically, we demonstrate new properties of the class of upper-$C^{1}$ functions [39], and in particular the convergence of Sequential Quadratic Programming (SQP) [29], a standard constrained optimization algorithm, when applied to such functions even though they are not differentiable everywhere. We deduce that SQP can be applied to the nonsmooth problem (3), and to (4) under an additional assumption, to find their Karush-Kuhn-Tucker points. From a practical point of view, we also study the use of Monte-Carlo sampling to perform a global exploration of the parameter space, with SQP subsequently ensuring efficient local exploitation. Then, following the approach of [3], this worst-case search method is combined with multi-models controller optimization [8] to perform robust control. The proposed approach is illustrated through the control of a mechanical system, which typically exhibits constraints of the form (5). To the best of our knowledge, this is the first robust $\mathcal{H}_{\infty}$ control approach that tackles nonlinear constraints on the uncertain parameters.
@@
We emphasize that the proposed approach, although studied here in the context of the constrained space (5), may also prove valuable in the unconstrained case. Indeed, $\mu$-analysis remains difficult to apply to industrial systems with many uncertain parameters. By contrast, the proposed approach scales well to plants with many states and parameters (as a gradient-based method), albeit without formal guarantee of global optimality – this problem being known to be NP-hard. Furthermore, the nonsmooth optimization of [3], implemented in systune, is local and sometimes misses worst-case configurations, whereas the proposed approach performs a global exploration of the parametric space.
@@ II-A Properties of upper-$C^{1}$ functions
[3, Sections III.A., II.B., V.A.] shows that minimization of such upper-$C^{1}$ functions, even when not differentiable everywhere, exhibits similar behavior to smooth functions. In particular, the authors show the convergence of their bundle method, in the search of the worst-case $\mathcal{H}_{\infty}$ norm, even when the cutting plane step is reduced to a standard linesearch (as in smooth optimization). This work follows [25] that treats of the optimization of lower- and upper-$C^{1}$ functions. Similar considerations are also thoroughly discussed by the same authors in [11]. Similarly, in [26], the Frank-Wolfe algorithm, originally designed for smooth functions, is also proven to converge to Clarke-stationary points when applied to upper-$C^{1}$ functions. In the same spirit, we now show that the upper-$C^{1}$ functions, although not differentiable everywhere, maintain some favorable properties for their minimization. Then, we will demonstrate the convergence of one particular smooth optimization algorithm when applied to upper-$C^{1}$ functions.
@@
We recall that the directional derivative at $x$ in direction $d$ of $f$ defined as
@@
always exists for an upper-$C^{1}$ function [35]. The subdifferential of $f$ at $x$ is denoted $\partial f(x)$.
@@
Let a locally Lipschitz function $f$ be upper-$C^{1}$ at $x$.
@@
b) Optimality. If $x$ locally minimizes $f(x)$, then $g_{x}||=0$ for any $g_{x}\in\partial f(x)$.
@@
a) [26, Proposition 2 and equation (6)] shows that the directional derivative of an upper-$C^{1}$ function $f$ is:
$f^{\prime}(x,d)=\min_{g\in\partial f(x)}g^{T}d=\min_{u\in I(x)}\nabla_{x}F(x,u)^{T}d$ (7)
@@
Let a locally Lipschitz function $f$ be upper-$C^{1}$ at $x$. Let the $c_{i}$ be continuously differentiable.
a) Descent direction. Let $(p,u)$ be a KKT pair of $Q(x,H,g_{x})$ with $g_{x}\in\partial f(x)$ any subgradient. If $p||\neq 0$ and $u||_{\infty}\leq r$, then $p$ is a descent direction for $\theta_{r}(x)$, i.e. $\theta^{\prime}_{r}(x,p)<0$.
b) Optimality. If $x$ locally minimizes $f(x)$ subject to $n$ constraints $c_{i}(x)\leq 0$ (Problem (8)), then any subgradient $g_{x}\in\partial f(x)$ verifies the following KKT conditions: there exists $(\mu_{i})_{1\leq i\leq n}$ such that:
@@
a) Cf. Appendix. The proof is a mere adaptation of [28, Theorem 3.1] to the upper-$C^{1}$ case.
@@
Proposition 1.b) is also proved in [26, Proposition 4]. Proposition 2.a) is a generalization of [28, Theorem 3.1], originally for differentiable functions, to upper-$C^{1}$ functions. To the best of our knowledge, Proposition 2.b) is novel, and can also be seen as a specific case of the usual KKT conditions for upper-$C^{1}$ functions. It allows optimality to be checked with any subgradient, in contrast to the general nonsmooth case, where only the existence of one subgradient satisfying KKT conditions is required.
@@ II-B Convergence of sequential quadratic programming (SQP) for upper-$C^{1}$ functions
We now introduce the Sequential Quadratic Programming (SQP) algorithm to solve Problem (8). It is described in Algorithm 1, as in the original reference [28] for differentiable functions, except that we replaced the gradient $\nabla f(x)$ by any subgradient $g_{x}\in\partial f(x)$ in the quadratic problem $Q(x,H,g_{x})$ (Eq. (10)). See also [33] for a discussion when the penalty parameter $r$ is variable. We do not detail how the quadratic problem in Step 1 or the step size $\lambda_{k}$ in Step 2 may be solved, or how $H_{k}$ is updated in Step 3, as it does not affect the convergence proof (as long as the conditions described in Algorithm 1 are satisfied)
@@
Step 1: Pick a $g_{x_{k}}\in\partial f(x_{k})$. Find a KKT point $p_{k}$ of the quadratic problem $Q(x_{k},H_{k},g_{x_{k}})$ defined in Eq. (10).
@@
where $\theta_{r}$ is defined in Eq. (11), and ($\epsilon_{k}$) is a sequence of nonnegative numbers satisfying:
@@
Proposition 3 extends the convergence result of SQP established in [28, Theorem 3.2] for differentiable objective functions to the case of upper-$C^{1}$ objective functions. Besides the upper-$C^{1}$ condition of $f$, the same assumptions as in [28, Theorem 3.2] are considered.
@@
(ii) For each $k$, there exists a KKT point of $Q(x_{k},H_{k},g_{x_{k}})$ (Eq. (10)) with a Lagrange multiplier vector $u_{k}$ such that $u_{k}||_{\infty}\leq r$ (vector $\infty$-norm).
@@
Cf. Appendix. The proof follows that of [28, Theorem 3.2] and adapts it to the case where f is not differentiable but only upper-$C^{1}$. In particular, the upper semi-continuity of the Clarke subdifferential allows to maintain the argument linked to the continuity of the gradient in the original proof. The existence of an accumulation point of the solutions $p_{k}$ of the quadratic problems $Q(x_{k},H_{k},g_{x_{k}})$ also follows from the submonotonicity of the subdifferential of lower-$C^{1}$ functions [39] (cf. Lemma 1 in Appendix). ∎
@@
The purpose of our arguments for using smooth optimization is not to claim superiority over specialized nonsmooth optimization methods (for example [25, 24] for lower- or upper-$C^{1}$ functions specifically). Instead, we emphasize that readily available algorithms (e.g., in Matlab or Python toolboxes) can be applied successfully.